Pfender Guitars has a current annual cash dividend policy of ​$4.00. The price of the stock is set to yield a return of 8​%. What is the price of this stock if the dividend will be paid:

a.  for 8 years and then the company repurchases the stock for $35​?

b.  for 11 years and then the company repurchases the stock for $35​?

c.  for 50 years and then the company repurchases the stock for $35​?

d.  for 60 years and then the company repurchases the stock for $35?

e.  for 100 years and then the company repurchases the stock for $35?

f.   forever with no repurchase of the​ stock?
Answer #1

Ans :

Price = Dividend × (1 – 1/(1 + r)n) / r + Liquidating Dividend × (1/(1 + r)n)

(A) $4.00 × (1 – 1/(1.08)8/ 0.08 + $35.00 × 1/(1.08)8 = 22.9865+18.099 = 41.895

(b) $4.00 × (1 – 1/(1.08)11/ 0.08 + $35.00 × 1/(1.08)11 = 28.55 + 15.010 = 43.566

(c) $4.00 × (1 – 1/(1.08)50/ 0.08 + $35.00 × 1/(1.08)50 = 48.933 + 0.7462 = 49.6801

(d) $4.00 × (1 – 1/(1.08)60/ 0.08 + $35.00 × 1/(1.08)60 = 49.5062 + 0.3456 = 49.8518

(e) $4.00 × (1 – 1/(1.08)100/ 0.08 + $35.00 × 1/(1.08)100 = 49.977 + 0.0159 = 49.993

(f) $4.00 /0.08 = 50